OpenStudy (anonymous):

what is the value of these expression 3/3^-2 k-4=1/256 2n=1/8

OpenStudy (ipwnbunnies):

Your 2nd question is similar to the first question: \[\frac{3}{3^(-2)}\]

OpenStudy (ipwnbunnies):

Whoops, that's supposed to be a power. But yeah, what did I say about negative exponents?

OpenStudy (anonymous):

would i do -3 *-3 than divied

OpenStudy (ipwnbunnies):

What? No. First step. Let's bring what is in the denominator to the numerator.

OpenStudy (anonymous):

?

OpenStudy (ipwnbunnies):

What did I just tell you. If you have a negative exponent, you can write it on the other side of the fraction and the exponent comes positive...I told you 3 times.

OpenStudy (ipwnbunnies):

3/3^(-2) = 3*3^(2) =

OpenStudy (anonymous):

would it be 18 ?